# Reflections on Diversity: A Real-time Virtual Mirror for Inclusive 3D Face Transformations

### v1

@@ Abstract
Real-time 3D face manipulation has significant applications in virtual reality, social media and human-computer interaction. This paper introduces a novel system, which we call Mirror of Diversity (MOD), that combines Generative Adversarial Networks (GANs) for texture manipulation and 3D Morphable Models (3DMMs) for facial geometry to achieve realistic face transformations that reflect various demographic characteristics, emphasizing the beauty of diversity and the universality of human features. As participants sit in front of a computer monitor with a camera positioned above, their facial characteristics are captured in real time. Our system provides a dynamic, responsive “mirror” effect, allowing the digital 3D model to follow the participant’s motions, offering an immersive virtual reflection. Participants can further alter their digital face reconstruction with transformations reflecting different demographic characteristics—such as gender and ethnicity (e.g., a person from Africa, Asia, Europe). Another feature of our system, which we call “Collective Face”, generates an averaged face representation from multiple participants’ facial data. As each new face is processed, identity coefficients and textures are averaged to continuously update this collective face. A comprehensive evaluation protocol is implemented to assess the realism and demographic accuracy of the transformations. Qualitative feedback is gathered through participant questionnaires, which include comparisons of MOD transformations with similar filters on platforms like Snapchat and TikTok, focusing on realism, feature preservation, and faithfulness to demographic representation. Additionally, quantitative analysis is conducted using a pretrained Convolutional Neural Network that predicts gender and ethnicity, to validate the accuracy of demographic transformations. Project webpage: https://vivianval.github.io/ReflectionsOnDiversity
@@ I Introduction
Our key contributions can be summarized as follows:
  • Development of the Mirror of Diversity (MOD) system, as part of the STEAMDIVE-Erasmus+ project, which allows users to experience realistic transformations of their faces as if they were of a different gender or ethnicity.
  • Introduction of a “Collective Face” feature that continuously aggregates facial data from participants, underscoring the shared humanity across diverse individuals.
  • Use of state-of-the-art methods like Large Scale Facial models to create highly realistic face transformations.
  • Engagement through a virtual interactive mirror.
  • Educational impact, since this tool can be used to facilitate conversations around identity, culture, and inclusiveness, allowing students to see themselves and others in various ethnic representations and fostering a more inclusive worldview.
@@ III-E Implementation
Our system was tested on an AMD Ryzen 9 7900 (12-core, 24-thread) processor, 61 GiB RAM, and NVIDIA RTX 3060 for GPU-accelerated processing.
@@ Acknowledgments
This work was co-funded by the project Innovative Educational Tools for Promoting Inclusion and Diversity in Schools (DIVERSITY IN STEAM) – STEAMDIVE, supported by the European Commission under the ERASMUS+ Programme (Cooperation Partnerships in School Education). For more information, visit: https://steamdive.eu.

### v2

@@ Abstract
Real-time 3D face manipulation has significant applications in virtual reality, social media and human-computer interaction. This paper introduces a novel system, which we call Mirror of Diversity (MOD), that combines Generative Adversarial Networks (GANs) for texture manipulation and 3D Morphable Models (3DMMs) for facial geometry to achieve realistic face transformations that reflect various demographic characteristics, emphasizing the beauty of diversity and the universality of human features. As participants sit in front of a computer monitor with a camera positioned above, their facial characteristics are captured in real time. Our system provides a dynamic, responsive “mirror” effect, allowing the digital 3D model to follow the participant’s motions, offering an immersive virtual reflection. Participants can further alter their digital face reconstruction with transformations reflecting different demographic characteristics—such as gender and ethnicity (e.g., a person from Africa, Asia, Europe). Another feature of our system, which we call “Collective Face”, generates an averaged face representation from multiple participants’ facial data. A comprehensive evaluation protocol is implemented to assess the realism and demographic accuracy of the transformations. Qualitative feedback is gathered through participant questionnaires, which include comparisons of MOD transformations with similar filters on platforms like Snapchat and TikTok, focusing on realism, feature preservation, and faithfulness to demographic representation. Additionally, quantitative analysis is conducted using a pretrained Convolutional Neural Network that predicts gender and ethnicity, to validate the accuracy of demographic transformations. Project webpage: https://vivianval.github.io/ReflectionsOnDiversity
@@ I Introduction
In this paper, we present the Mirror of Diversity (MOD) computer vision software application, as part of the STEAMDIVE project, which allows users to experience realistic transformations of their faces as if they were of a different gender or ethnicity. This educational tool is in the form of a virtual interactive mirror that raises awareness around identity, multiculturalism, promoting inclusion and diversity in an educational environment.
@@ II Related Work
The concept of virtual mirror has been utilized in various behavioural [5], psychological [6] and healthcare applications [7, 8].
@@ III-E Implementation
Our system was tested on an AMD Ryzen 9 7900 (12-core, 24-thread) processor, 61 GiB RAM, and NVIDIA RTX 3060 for GPU-accelerated processing. The system maintains a frame rate of 60 FPS as reported by the OpenGL rendering loop.
@@ Acknowledgments
This work was co-funded by the project STEAMDIVE: Diversity in STEAM, Innovative Educational Tools for Promoting Inclusion and Diversity in Schools, supported by the European Commission under the ERASMUS+ Programme, KA220-SCH - Cooperation partnerships in school education (GA: 2022-1-EL01-KA220-SCH-000086968).
This work was also co-funded by the Hellenic Foundation for Research and Innovation (HFRI) under the “1st Call for HFRI Research Projects to support Faculty members and Researchers and the procurement of high-cost research equipment”, project I.C.Humans, no 91.